A satellite of mass 42.5 kg in geosynchronous orbit at an altitude of 3.58 ✕ 104 km above the Earth's surface remains above the same spot on the Earth. Assume its orbit is circular. Find the magnitude of the gravitational force exerted by the Earth on the satellite. Hint: The answer is not 417 N
Answer #1

Step:1

Given that

Mass of satellite m=42.5 kg

Distance h=3.58×10^4 km=3.58×10^7 m

radius of earth r=6400 km=0.64*10^7 m

Mass of earth M=5.972×10^24 kg

Step:2

Now we find the gravitational force

Force F=GMm/(r+h)^2

=6.67×10^-11×5.972×10^24×42.5/(0.64*10^7+3.58×10^7)^2

=95.1*10^-1

=9.51 N